OpenStudy (anonymous):

The height above ground,h, of an arrow t seconds after being shot straight in the air can be modeled by the formula h(t)=1.5+40t-4.9t^2, t= seconds h=meters. What is average rate of change of the height,in meters per second,of the first three second after being shot?

OpenStudy (amistre64):

\[\frac{h(3)-h(0)}{3-0}\]

OpenStudy (amistre64):

in general, the slope of the line between (0,h(0)) and (3,h(3))

OpenStudy (anonymous):

what do you do with the equation?

OpenStudy (phi):

you are given h(t)=1.5+40t-4.9t^2 to find h(3), erase the t in the equation, and put in 3 instead. then do the arithmetic for h(0), do the same thing, except use 0. example: h(0)= 1.5 + 40*0 - 4.9 * 0^2 if you simplify you get h(0)= 1.5+ 0 - 0 = 1.5 so h(0)= 1.5 can you find h(3) ?

OpenStudy (anonymous):

yes thank you